Determine analytically if the following functions are even, odd, or neither
a. f(x) = 5x ^3/(2x) b. f(x) = x ^2 + 3x + 2

Answer by CubeyThePenguin(3113)

a) f(x) = 2x^2
f(-x) = 2(-x)^2 = 2x^2

f(x) is even.

b) f(x) = x^2 + 3x + 2
f(-x) = (-x)^2 + 3(-x) = 2 = x^2 - 3x + 2

f(x) is neither even nor odd.